Find the roots of 2x^4 + 9x^3 -9x^2 - 46x +24 by Ferrari’s method.
1
Expert's answer
2020-02-26T11:36:31-0500

To solve this equation by Ferrari’s method, first divide each term of the equation by the leading coefficient (2). Obtain


"x^4 + \frac{9}{2}x^3 - \frac{9}{2}x^2 - 23x + 12 = 0"


Denote the coefficients of the obtained equation as "a=\frac{9}{2}, b=-\frac{9}{2}, c=-23, d=12".

Next, you should determine any root of the following equation.


"y^{3}-by^{2}+(ac-4d)y-a^{2}d+4bd-c^{2}=0"


Substituting "a,b,c,d", obtain


"y^3 + \frac{9}{2}y^2 -\frac{300}{2}y-988=0"


By substitution method determine that the root is "y_0 = -8".

Then the needed roots is the roots of the next equations.


"x^{2}+{\frac{a}{2}}x+{\frac {y_{0}}{2}}=\pm {\sqrt {\left({\frac{a^{2}}{4}}-b+y_{0}\right)x^{2}+\left({\frac{a}{2}}y_{0}-c\right)x+{\frac{y_{0}^{2}}{4}}-d}}"


Substituting "a,b,c,d, \text{and } y_0", obtain


"x^{2}+{\frac{9}{24}}x - 4=\pm \sqrt{\frac{25}{16}x^2 + 5x + 4}"


The square-root expressions is the perfect square.


"x^{2}+{\frac{9}{24}}x - 4=\pm \sqrt{\left(\frac{5}{4}x + 2\right)^2}"


Thus, obtain two equation. Solve the first equation.


"x^{2}+{\frac{9}{24}}x - 4=\frac{5}{x}x+2 \\ \\
x^{2}+x-6=0"


The roots are "-3" and "2".


Solve the second equation.


"x^{2}+{\frac{9}{24}}x - 4=-\frac{5}{x}x-2 \\ \\
x^{2}+\frac{7}{2}x-2=0"


The roots are "-4" and "\frac{1}{2}".


Therefore, the roots of the original equation are "-4, -3, \frac{1}{2}, \text{and } 2".